OpenStudy (anonymous):

Solve the system of equations. 3x – 7y = 5 9y = 5x + 5 A. (–3, 2) B. (8, 5) C. (–10, –5) D. (0, 0)

OpenStudy (whpalmer4):

\[3x-7y=5\]\[9y=5x+5\]Let's write the equations in the same order \[3x-7y=5\]\[-5x+9y=5\]To solve this by elimination, I would multiply the first equation by 5, and the second equation by 3. That will give you \(x\) terms of \(15x\) and \(-15x\), so if you then add the equations together straight down the columns, you'll end up with an equation only involving \(y\). You solve that for the value of \(y\), and then use that value in one of the equations to find the value of \(x\). Of course, with a multiple-guess question like this, where there is only 1 answer choice per value of \(y\), you don't really have to do the last step, do you? Still, it would be a good idea, you can probably use the practice, and when you're done, plug both numbers into both equations and make sure that the equations are true — if not, you've made a mistake somewhere.

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

@uri can u help?

OpenStudy (whpalmer4):

Did you test your solution in both equations? If you did what I told you, you'll know the answer, no need to ask if it is correct.

OpenStudy (whpalmer4):

But let's check: \[3x-7y=5\]You think that \((0,0)\) is the answer, so let's plug it in: \[3(0)-7(0) = 5\]\[0-0=5\]Is that true?

OpenStudy (whpalmer4):

How did you come to choose \((0,0)\) as your answer? If you show me your work, we can find the mistake and fix it.